# Second-Order Fine-Tuning without Pain for LLMs: A Hessian Informed Zeroth-Order Optimizer

### v1

@@ 3.2 Diagonal Hessian Estimator
3.3 Natural Gradient Estimation
  Input: parameters $\theta\in\mathbb{R}^{d}$, loss $L:\mathbb{R}^{d}\rightarrow\mathbb{R}$, step budget $T$, perturbation scale $\mu$, learning rate schedule ${\eta_{t}}$, smooth scale $\alpha_{t}$, diagonal Hessian $\Sigma_{0}$
  for $t=1,...,T$ do
     Sample batch $\mathcal{B}\subset\mathcal{D}$ and random seed $s$
     $\ell\leftarrow\mathcal{L}(\theta;\mathcal{B})$
     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
     $\ell_{+}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
     $\theta$leftarrow PerturbParameters($\theta$, $-2\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
     $\ell_{-}\leftarrow\mathcal{L}(\theta;\mathcal{B})$
     //Reset parameters before descent
     $\theta$leftarrow PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t-1}$, $s$)
     $\Sigma^{\prime}_{t}=\frac{1}{2\mu^{2}}(\ell_{+}+\ell_{-}-2\ell)(\Sigma^{-1/2}_{t-1}u_{i}u_{i}^{\top}\Sigma^{-1/2}_{t-1})$
     //Update diagonal Hessian
     $\Sigma_{t}=(1-\alpha_{t})\Sigma_{t-1}+\alpha_{t}\left|diag(\Sigma^{\prime}_{t})\right|$
     projected_grad $\leftarrow(\ell_{+}-\ell_{-})\ast\Sigma^{1/2}_{t}/2\mu$
     //For sampling $u_{i}$
     Reset random number generator with seed $s$
     Update parameters in place
     for $\theta_{i}\in\theta$ do
        Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
        $\theta_{i}\leftarrow\theta_{i}-\eta_{t}\ast$ projected_grad $\ast u_{i}$
     end for
  end for
  PerturbParameters($\theta$, $\mu$, $\Sigma^{1/2}_{t}$, $s$)
  //For sampling $u_{i}$
  Reset random number generator with seed $s$
  //Modify parameters in place
  for $\theta_{i}\in\theta$ do
     Sample $u_{i}\sim\mathcal{N}(0,I_{d})$
     $\theta_{i}\leftarrow\theta_{i}+\mu\Sigma^{1/2}_{t}u_{i}$
     Return $\theta$
  end for
Algorithm 1 HiZOO
Given the approximate Hessian inverse matrix $\Sigma$, we can construct the following descent direction:
$\begin{split}g_{\mu}(\theta)=\sum\limits_{i=1}^{n}\frac{\mathcal{L}(\theta+\mu\Sigma^{1/2}_{t}u_{i})-\mathcal{L}(\theta-\mu\Sigma^{1/2}_{t}u_{i})}{2\mu\cdot n}\cdot\Sigma_{t}^{1/2}u_{i}.\end{split}$ (5)
With the gradient descent direction, we can update $x_{t}$ as follows:
$\theta_{t+1}=\theta_{t}-\eta_{t}g_{\mu}(\theta_{t}).$ (6)
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1})\right]$ (7)
$=$ $\mathcal{L}(\theta_{t})-\eta_{t}\mathbb{E}\left[\langle\nabla\mathcal{L}(\theta_{t}),g_{\mu}(\theta_{t})\right]+\mathcal{O}(\eta_{t}^{2})$
$=$ $\mathcal{L}(\theta_{t})-\eta_{t}\frac{1}{b}\mathbb{E}\left[\sum_{i=1}^{b}\langle\nabla\mathcal{L}(\theta_{t}),\Sigma_{t}^{1/2}u_{i}u_{i}^{\top}\Sigma_{t}^{1/2}\nabla\mathcal{L}(\theta_{t})\right]$
$+\mathcal{O}(\eta_{t}^{2})+\mathcal{O}(\mu)$
$=$ $\mathcal{L}(\theta_{t})-\eta_{t}\|\Sigma_{t}^{1/2}\nabla(\theta_{t})\|^{2}+\mathcal{O}(\eta_{t}^{2})+\mathcal{\mu},$
4 Convergence Analysis
Task Type SST-2 SST-5 SNLI MNLI RTE TREC
—— sentiment —— —— natural language inference —— — topic —
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0
LP 76.0 ($\pm$2.8) 40.3 ($\pm$1.9) 66.0 ($\pm$2.7) 56.5 ($\pm$2.5) 59.4 ($\pm$5.3) 51.3 ($\pm$5.5)
FT 91.9 ($\pm$1.8) 47.5 ($\pm$1.9) 77.5 ($\pm$2.6) 70.0 ($\pm$2.3) 66.4 ($\pm$7.2) 85.0 ($\pm$2.5)
FT (LoRA) 91.4 ($\pm$1.7) 46.7 ($\pm$1.1) 74.9 ($\pm$4.3) 67.7 ($\pm$1.4) 66.1 ($\pm$3.5) 82.7 ($\pm$4.1)
FT (prefix) 91.9 ($\pm$1.0) 47.7 ($\pm$1.1) 77.2 ($\pm$1.3) 66.5 ($\pm$2.5) 66.6 ($\pm$2.0) 85.7 ($\pm$1.3)
MeZO 90.5 ($\pm$1.2) 45.5 ($\pm$2.0) 68.5 ($\pm$3.9) 58.7 ($\pm$2.5) 64.0 ($\pm$3.3) 76.9 ($\pm$2.7)
MeZO (LoRA) 91.4 ($\pm$0.9) 43.0 ($\pm$1.6) 69.7 ($\pm$6.0) 64.0 ($\pm$2.5) 64.9 ($\pm$3.6) 73.1 ($\pm$6.5)
MeZO (prefix) 90.8 ($\pm$1.7) 45.8 ($\pm$2.0) 71.6 ($\pm$2.5) 63.4 ($\pm$1.8) 65.4 ($\pm$3.9) 80.3 ($\pm$3.6)
HiZOO 93.2 ($\pm$0.8) 46.2 ($\pm$1.1) 74.6 ($\pm$1.3) 64.9 ($\pm$1.7) 66.8 ($\pm$ 1.2) 79.8 ($\pm$1.3)
HiZOO(LoRA) 91.2 ($\pm$0.4) 42.9 ($\pm$0.8) 71.1 ($\pm$1.1) 62.1 ($\pm$1.7) 61.4 ($\pm$2.2) 52.2 ($\pm$3.2)
HiZOO(prefix) 92.3 ($\pm$1.2) 47.2 ($\pm$1.1) 68.8 ($\pm$1.7) 61.6 ($\pm$1.2) 65.4 ($\pm$1.2) 82.0 ($\pm$2.0)
Table 1: Experiments on RoBERTa-large (350M parameters, k=16) that include zero-shot learning, linear probing (LP), full-parameter fine-tuning with Adam, MeZO and HiZOO, and parameter-efficient fine-tuning (LoRA and prefix tuning) with Adam, MeZO and HiZOO respectively. All reported numbers are averaged accuracy (standard deviation) across 5 runs. With full-parameter fine-tuning, HiZOO outperforms MeZO with 3.56% on average on all datasets.
Figure 3: HiZOO achieves a 5×5$\times$5 × speedup on average over MeZO. (a) HiZOO achieves 64.9% while MeZO achieves 63%. (b) HiZOO (LoRA) achieves 62.1% while MeZO (LoRA) achieves 61%. (c) HiZOO achieves 61.6% while MeZO achieves 59.9%. All experiments are conducted using the same parameters in Appendix D.1.
In Section 3, we have briefly discussed that $g_{\mu}(x)$ can estimate the gradient descent direction. In this section, we will analyse that our HiZOO will converge to a stationary point. Firstly, our convergence analysis requires the following assumptions:
Assumption 4.1.
The objective function $\mathcal{L}(\theta)$ is $L$-smooth, which means that for any $\theta_{1},\theta_{2}\in\mathbb{R}^{d}$, it holds that:
$\mathcal{L}(\theta_{2})\leq\mathcal{L}(\theta_{1})+\langle\nabla\mathcal{L}(\theta_{1}),\theta_{2}-\theta_{1}\rangle+\frac{L}{2}\|\theta_{2}-\theta_{1}\|^{2}.$ (8)
Assumption 4.2.
The stochastic gradient $\nabla\mathcal{L}(\theta,\mathcal{B})$ has $\sigma^{2}$ variance, which means:
$\mathbb{E}\left[\|\nabla\mathcal{L}(\theta,\mathcal{B})-\nabla\mathcal{L}(\theta)\|^{2}\right]\leq\sigma^{2}.$ (9)
Assumption 4.3.
Based on the above assumption of non-convex optimization, we can get the following convergence analysis.
@@ 4 Convergence Analysis
Let the approximate gradient $g_{\mu}(\theta_{t})$ defined as:
$g_{\mu}(\theta_{t})=\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i})}{2b\mu}\Sigma_{t}^{1/2}u_{i}.$ (10)
Based on Assumption 4.1-4.3, we can draw the following corollary. If the update rule for $\theta$ is $\theta_{t+1}=\theta_{t}-\eta g_{\mu}(\theta_{t})$ for a single step, then it’s established that:
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1})\mid\right]\leq\mathcal{L}(\theta_{t})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t})\|_{\Sigma_{t}}^{2}$ (11)
$+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}).$
[…]
where $\mathcal{L}(\theta_{*})$ minimizes the function $\mathcal{L}(\theta)$. The above equation shows that as $T\to\infty$, HiZOO can converge to the optimum point.
@@ 5 Experiments
Large language models are generally classified into two types: (1) Encoder-Decoder, also known as masked language models, such as BERT (Devlin et al., 2019) and ALBERT (Lan et al., 2020); (2) Decoder-Only, also recognized as generative language models, such as GPT family (Brown et al., 2020), OPT (Zhang et al., 2022a), LLaMA (Touvron et al., 2023).
Task SST-2 RTE CB BoolQ WSC WIC MultiRC COPA ReCoRD SQuAD
Task type ———————— classification ———————— — multiple choice — generation
Zero-shot 58.8 59.6 46.4 59.0 38.5 55.0 46.9 80.0 81.2 46.2
ICL 87.0 62.1 57.1 66.9 39.4 50.5 53.1 87.0 82.5 75.9
LP 93.4 68.6 67.9 59.3 63.5 60.2 63.5 55.0 27.1 3.7
FT (6x memory) 92.0 70.8 83.9 77.1 63.5 70.1 71.1 79.0 74.1 84.9
MeZO 91.4 66.1 67.9 67.6 63.5 61.1 60.1 88.0 81.7 84.7
MeZO(LoRA) 89.6 67.9 66.1 73.8 64.4 59.7 61.5 84.0 81.4 83.8
MeZO(prefix) 90.7 70.8 69.6 73.1 57.7 59.9 63.7 87.0 81.2 84.2
HiZOO 91.3 69.3 69.4 67.3 63.5 59.4 55.5 88.0 81.4 81.9
HiZOO(LoRA) 90.59 67.5 69.6 70.5 63.5 60.2 60.2 87.0 81.9 83.8
HiZOO(prefix) 92.0 71.8 69.6 73.9 60.6 60.0 64.8 87.0 81.2 83.2
Table 2: Experiments on OPT-13B (with $1000$ examples). ICL: in-context learning; LP: linear probing; FT: full fine-tuning with Adam. We highlight the best results between HiZOO and MeZO in bold to facilitate comparison. HiZOO outperforms zero-shot, ICL, and LP across the board, and achieves comparable (within 1%) or better performance.
Figure 4: HiZOO greatly accelerates the speed of model convergence over MeZO. (a) HiZOO (LoRA) achieves 67.5% while MeZO (LoRA) achieves 67.9% finally on RTE dataset. (b) HiZOO (prefix) achieves 69.6% while MeZO (LoRA) achieves 69.6% finally on CB dataset. (c) HiZOO (LoRA) achieves 60.6% while MeZO (LoRA) achieves 57.7% finally on WSC dataset. Detailed hyper-parameters can be found in Appendix E.
To rigorously assess the universality and robustness of our HiZOO, we have chosen models from each category for empirical testing. Additionally, we investigate full-parameter tuning and PEFT methods, which encompasses methods such as Low-Rank Adaptation (LoRA) (Hu et al., 2022) and prefix-tuning (Li & Liang, 2021). Detailed methodologies are presented in Appendix D.2.
@@ 5.1 Masked Language Models
Firstly, we conduct experiments on RoBERTa-large 350M (Liu et al., 2019b) on three NLP task paradigms: sentence classification, multiple choice and text generation. We follow the experimental setting (Malladi et al., 2023) in studying the few-shot and many-shot, sampling $k$ examples per class for $k=16$ (results in Table 1) and $k=512$ (results in Appendix D.2), which can reach the following observations and summaries.
HiZOO greatly increases the convergence speed across full-parameter tuning, LoRA and prefix. As shown in Figure 3, HiZOO achieves 5× speedup over MeZO on average while getting the same training loss compared with MeZO. What’s more, the acceleration effect is stable across three tuning methods. More figures for curves of training loss are shown in Appendix D.2.
HiZOO achieves better performance compared with MeZO. Table 1 shows that HiZOO outperforms MeZO’s results with 3.56% on average on all datasets across sentiment, natural language inference and topic tasks. Specifically, HiZOO outperforms MeZO with more than 6% on SNLI and MNLI dataset.
Task SST-2 RTE BoolQ WSC WIC
30B zero-shot 56.7 52.0 39.1 38.5 50.2
30B ICL 81.9 66.8 66.2 56.7 51.3
30B MeZO 90.6 66.4 67.2 63.5 56.3
30B MeZO(prefix) 87.5 72.6 73.5 55.8 59.1
30B HiZOO 90.3 69.3 70.8 63.5 53.4
30B HiZOO(prefix) 91.2 68.6 73.1 57.7 60.2
66B zero-shot 57.5 67.2 66.8 43.3 50.6
66B ICL 89.3 65.3 62.8 52.9 54.9
66B MeZO(prefix) 93.6 66.4 73.7 57.7 58.6
66B HiZOO(prefix) 93.6 71.5 73.2 60.6 61.1
Table 3: Experiments on OPT-30B and OPT-66B (with $1000$ examples). The best results of the two methods are both highlighted in bold for better comparison. We highlight the best results between HiZOO and MeZO in bold to facilitate comparison. We see that on most datasets HiZOO outperforms MeZO.
@@ 5.2 Auto-Regressive Language Models
Then we extend experiments with OPT family on the same NLP task paradigms. The main results on OPT-13B are presented in Table 2. We can see that HiZOO achieves comparable (within 1%) or better performance than MeZO on all datasets. What’s more, HiZOO works well on both full-parameter tuning and PEFT methods.
HiZOO is capable of scaling to large models with up to 66B parameters, while preserving its exceptional performance. As depicted in Table 3, on OPT-66B HiZOO (prefix) outperforms MeZO(prefix) with up to 5% increase and 2% increase on average. Figure 4 shows that at such scales HiZOO can still effectively accelerate the process for convergence and reduce the training steps.

### v3

@@ 3.3 Diagonal Hessian Estimator
To further reduce Hessian memory consumption, we propose HiZOO-L to maintain it in a low-rank subspace, motivated by Adafactor [41]. For $\hat{\Sigma}^{-1}\in\mathbb{R}^{p\times q}$, we will store two low-rank matrices $R\in\mathbb{R}^{p\times k}$ and $C\in\mathbb{R}^{k\times q}$ with $k=1$. Specifically, we can get $\hat{\Sigma}^{-1}$ by:
$\hat{\Sigma}^{-1}_{t}=(R_{t}*C_{t})/(1_{p}^{\top}*R_{t}),$
where $1_{p}=(1,\cdots,1)\in\mathbb{R}^{p}$ denotes a column vector of $p$ ones. Then in each step, we will update the $R$ and $C$ separately:
$R_{t}^{-1}=(1-\alpha_{t})R_{t-1}^{-1}+\alpha_{t}\left|diag(\hat{\Sigma}^{\prime}_{t})\right|*1_{q},$
$C_{t}^{-1}=(1-\alpha_{t})C_{t-1}^{-1}+\alpha_{t}1^{\top}_{p}*\left|diag(\hat{\Sigma}^{\prime}_{t})\right|.$
Detailed Algorithm can be seen in Appendix D.
3.4 Convergence Analysis
In this section, we will analyse the convergence based on the assumption of non-convex optimization (details in Appendix B).
@@ 3.4 Convergence Analysis
Let the descent direction $g_{\mu}(\theta_{t})$ defined as:
$g_{\mu}(\theta_{t})=\sum_{i=1}^{b}\frac{\mathcal{L}(\theta_{t}+\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B})-\mathcal{L}(\theta_{t}-\mu\Sigma_{t}^{1/2}u_{i};\mathcal{B})}{2b\mu}\Sigma_{t}^{1/2}u_{i}.$ (7)
Based on Assumption B.1-B.3, if the update rule for $\theta$ is $\theta_{t+1}=\theta_{t}-\eta g_{\mu}(\theta_{t})$ for a single step, then it’s established that:
$\mathbb{E}\left[\mathcal{L}(\theta_{t+1};\mathcal{B})\mid\right]\leq\mathcal{L}(\theta_{t};\mathcal{B})-\frac{\eta_{t}}{4}\|\nabla\mathcal{L}(\theta_{t};\mathcal{B})\|_{\Sigma_{t}}^{2}+2\eta_{t}^{2}L\left(\mathrm{tr}(\Sigma_{t})+\beta_{u}\right)\sigma^{2}+\mathcal{O}(\mu^{2}).$ (8)
Furthermore, given iteration number $T$, we choose the step size $\eta=\frac{1}{8\sqrt{T}L(\max_{t}\mathrm{tr}(\Sigma_{t})+\beta_{u})}$ and take $\theta_{\mbox{out}}=\theta_{j}$ with $j$ uniformly sampled from $\{1,\dots,T\}$. Then, we have
$\begin{split}\mathbb{E}\left[\|\nabla\mathcal{L}(\theta_{\mbox{out}};\mathcal{B})\|^{2}\right]\leq\frac{32L\left(\max_{t}\{\mathrm{tr}(\Sigma_{t})\}+\beta_{u}\right)(\mathcal{L}(\theta_{1};\mathcal{B})-\mathcal{L}(\theta_{*};\mathcal{B}))}{\sqrt{T}\beta_{\ell}}+\frac{\sigma^{2}}{T^{3/2}\beta_{\ell}}+\mathcal{O}\left(\mu^{2}\right),\end{split}$ (9)
where $\mathcal{L}(\theta_{*};\mathcal{B})$ minimizes the function $\mathcal{L}(\theta;\mathcal{B})$. The above equation shows that as $T\to\infty$, HiZOO can converge to the optimum point.
@@
3.5 Visualization of HiZOO on Test Functions
Despite above theoretical guarantee, we still want to illustrate how HiZOO utilizes Hessian to improve the convergence process. But it’s impractical for large models to visualize their optimization trajectories. Therefore we choose three test functions (see details in Appendix C) with heterogeneous curvatures across different parameters and visualize the optimization trajectories on them.
As illustrated in Figure 3, HiZOO and Adam both achieve better convergence on three functions, and HiZOO even requires less steps for convergence than Adam. However, MeZO only achieves effective convergence in either the $x$ or $y$ dimension, but not both, indicating a limitation in capturing this curvature difference. Particularly in function (c) curvature of $x$ is extremely bigger than $y$. In this case, HiZOO can sense this difference in parametric curvature and update the function along $x$ on purpose, achieving quicker convergence. In contrast, MeZO is very hard to converge.
Figure 3: Optimization trajectories of Adam, MeZO and HiZOO on 3 test functions. We have labeled the number of iterations required for the loss to drop to 0.1.
@@ 4 Experiments
Large language models are generally classified into two types: (1) Encoder-Decoder, also known as masked language models, such as BERT [11] and ALBERT [23]; (2) Decoder-Only, also recognized as generative language models, such as GPT family [37, 3], OPT [57], LLaMA [45], Phi [25, 17].
To rigorously assess the universality and robustness of our HiZOO, we have chosen models from each category for empirical testing. Additionally, we investigate FT and PEFT (LoRA [19] and prefix [24]). Detailed experiment settings are presented in Appendix E.1.
@@ 4.1 Masked Language Models
Firstly, we conduct experiments on RoBERTa-large 350M [28] on three NLP task paradigms: sentence classification, multiple choice and text generation. We follow the experimental setting [32] in studying the few-shot and many-shot, sampling $k$ examples per class for $k=16$ (results in Table 1), which can reach the following observations and summaries.
Task Type SST-2 SST-5 SNLI MNLI RTE TREC Average
—— sentiment —— —— natural language inference —— — topic —
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0 49.5
LP 76.0 ($\pm$2.8) 40.3 ($\pm$1.9) 66.0 ($\pm$2.7) 56.5 ($\pm$2.5) 59.4 ($\pm$5.3) 51.3 ($\pm$5.5) 58.3
FT 91.9 ($\pm$1.8) 47.5 ($\pm$1.9) 77.5 ($\pm$2.6) 70.0 ($\pm$2.3) 66.4 ($\pm$7.2) 85.0 ($\pm$2.5) 74.9
PEFT 91.9 ($\pm$1.0) 47.7 ($\pm$1.1) 77.2 ($\pm$1.3) 67.7 ($\pm$1.4) 66.6 ($\pm$2.0) 85.7 ($\pm$1.3) 72.8
MeZO 90.5 ($\pm$1.2) 45.5 ($\pm$2.0) 68.5 ($\pm$3.9) 58.7 ($\pm$2.5) 64.0 ($\pm$3.3) 76.9 ($\pm$2.7) 67.4
MeZO (PEFT) 91.4 ($\pm$0.9) 45.8 ($\pm$2.0) 71.6 ($\pm$2.5) 64.0 ($\pm$2.5) 65.4 ($\pm$3.9) 80.3 ($\pm$3.6) 69.8
HiZOO 93.2 ($\pm$0.8) 46.2 ($\pm$1.1) 74.6 ($\pm$1.3) 64.9 ($\pm$1.7) 66.8 ($\pm$1.2) 79.8 ($\pm$1.3) 70.9
HiZOO(PEFT) 92.3 ($\pm$1.2) 47.2 ($\pm$1.1) 71.1 ($\pm$1.1) 62.1 ($\pm$1.7) 65.4 ($\pm$1.2) 82.0 ($\pm$2.0) 70.0
Table 1: Experiments on RoBERTa-large (350M parameters, k=16). PEFT represents the LoRA and prefix and we report the best of them. All reported numbers are averaged accuracy (standard deviation) across 5 runs.
Figure 4: OPT-13B results with zero-shot, in-context learning (ICL), MeZO, HiZOO and fine-tuning with Adam (FT). PEFT represents the best among LoRA and prefix. More details can be found in Appendix E.2.
HiZOO greatly increases the convergence speed across full-parameter tuning, LoRA and prefix. As shown in Figure 4, HiZOO achieves 4×4$\times$4 × speedup over MeZO on average while getting the same training loss compared with MeZO. What’s more, HiZOO finally achieves a 1.57% accuracy improvement better than MeZO. More loss curves can be seen in Appendix E.1.
HiZOO achieves better performance compared with MeZO. Table 1 shows that HiZOO outperforms MeZO’s results with 3.5% on average on all datasets across different tasks. Specifically, HiZOO outperforms MeZO more than 6% in accuon SNLI and MNLI dataset.
@@ 4.2 Auto-Regressive Language Models
Then we extend experiments with OPT family (13B, 30B, 66B) on the same NLP task paradigms. The main results on OPT-13B are presented in Table 2. We can see that HiZOO boosts MeZO’s accuracy for 1.1% accuracy on average using prefix-tuning. Specifically, HiZOO achieves 2.87% and 3% accuracy improvement on WSC and COPA tasks.
HiZOO is capable of scaling to large models with up to 66B parameters, while preserving its exceptional performance. As depicted in Table 4, on OPT-30B HiZOO outperforms MeZO with up to 2.9% increase and 0.8% increase on average. Even scaling to OPT-66B, HiZOO(prefix) still outperforms MeZO(prefix) with up to 5.1% increase and 2% increase on average.
In addition, We utilize HiZOO to fine-tune the newly proposed model, Phi-2 (2.7B) and Llama3 (8B). Experiment results in Table 2 show that HiZOO outperforms MeZO in most cases. We also provide relative loss curves to show the better convergence process of our HiZOO in Appendix E.3.
Table 2: Experiments on three different models(with $1000$ examples). We highlight the best results between MeZO, HiZOO and HiZOO-L in bold to facilitate comparison.
Model Method SST-2 RTE  CB WSC WIC COPA MultiRC Average
Phi-2 MeZO 86.6 67.1 75.0 59.6 54.4 86.0 78.2 72.4
Phi-2 HiZOO 88.9 69.0 75.2 62.5 59.4 86.0 79.2 74.3
Phi-2 HiZOO-L 88.9 68.9 75.2 62.4 59.2 86.0 79.2 74.2
Llama3 MeZO 92.2 74.4 69.6 63.5 57.8 88.0 77.6 74.7
Llama3 HiZOO 93.5 75.1 69.6 63.5 59.7 89.0 78.2 75.5
Llama3 HiZOO-L 94.3 75.1 69.6 63.5 57.7 89.0 77.9 75.3
OPT-13B MeZO 91.4 66.1 66.0 63.5 59.4 88.0 57.3 70.2
OPT-13B HiZOO 92.1 69.3 69.6 63.5 59.4 89.0 61.3 72.1
OPT-13B HiZOO-L 92.1 68.2 67.9 65.4 59.4 89.0 61.1 71.9